# Best approximation of a three-variable function by sum of one-variable coordinate functions

### v1

@@ 3. Case 3:
For any set $S=S_{x}\times S_{y}\times S_{z}$ where, $S_{x},S_{y},S_{z}\subset[0,1]$ are some finite sets with $\{0,1\}\subset S_{x}\cap S_{y}\cap S_{z}$ and for any function $f\in C(\Omega)$ satisfying the inequalities (2.1), the intersection $X^{\prime}_{S}\cap X_{S,f}$ is non-empty.
@@
$f_{n}(x,y,z)=f(x,y,z)+\frac{1}{n}(xy+xz+yz),\forall n\in\mathbb{N}$
@@
$\max_{(p,\lambda)\in X_{S}}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}=\max_{(p,\lambda)\in X_{S,min}}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}=\max_{(p,\lambda)\in X_{S,min}}\frac{\sum\lambda_{i}(\lim_{n\rightarrow\infty}f_{n}(x_{i}))}{\sum|\lambda_{i}|}$
Here, we can change the order of the limit and the maximum as $X_{S,min}$ is a finite set by the Lemma 2.3 (considering vectors up to multiplication by a constant) and hence,
$\max_{(p,\lambda)\in X_{S,min}}\frac{\sum\lambda_{i}(\lim_{n\rightarrow\infty}f_{n}(x_{i}))}{\sum|\lambda_{i}|}=\lim_{n\rightarrow\infty}\max_{(p,\lambda)\in X_{S,min}}\frac{\sum\lambda_{i}f_{n}(x_{i})}{\sum|\lambda_{i}|}=\lim_{n\rightarrow\infty}\max_{(p,\lambda)\in X_{S}}\frac{\sum\lambda_{i}f_{n}(x_{i})}{\sum|\lambda_{i}|}$
@@
$\lim_{n\rightarrow\infty}\max_{(p,\lambda)\in X_{S}}\frac{\sum\lambda_{i}f_{n}(x_{i})}{\sum|\lambda_{i}|}=\lim_{n\to\infty}\max_{(p,\lambda)\in X^{\prime}_{S}}\frac{\sum\lambda_{i}f_{n}(x_{i})}{\sum|\lambda_{i}|}=\lim_{n\to\infty}\max_{(p,\lambda)\in X^{\prime}_{S,min}}\frac{\sum\lambda_{i}f_{n}(x_{i})}{\sum|\lambda_{i}|}$
@@
$\lim_{n\rightarrow\infty}\max_{(p,\lambda)\in X^{\prime}_{S,min}}\frac{\sum\lambda_{i}f_{n}(x_{i})}{\sum|\lambda_{i}|}=\max_{(p,\lambda)\in X^{\prime}_{S,min}}\frac{\sum\lambda_{i}(\lim_{n\rightarrow\infty}f_{n}(x_{i}))}{\sum|\lambda_{i}|}=\max_{(p,\lambda)\in X^{\prime}_{S}}\frac{\sum\lambda_{i}(f(x_{i}))}{\sum|\lambda_{i}|}$
Now let’s choose $(p^{0},\lambda^{0})\in X^{\prime}_{S}$ such that,
$\frac{\sum\lambda^{0}_{i}(f(x_{i}))}{\sum|\lambda^{0}_{i}|}=\max_{(p,\lambda)\in X^{\prime}_{S}}\frac{\sum\lambda_{i}(f(x_{i}))}{\sum|\lambda_{i}|}.$
@@
$\frac{\sum\lambda^{0}_{i}(f(x_{i}))}{\sum|\lambda^{0}_{i}|}=\max_{(p,\lambda)\in X_{S}}\frac{\sum\lambda_{i}(f(x_{i}))}{\sum|\lambda_{i}|}$
and thus, $(p^{0},\lambda^{0})\in X_{S,f}$ by the definition of $X_{S,f}$. Hence, $(p^{0},\lambda^{0})\in X^{\prime}_{S}\cap X_{S,f}$
@@ 2.2 Construction of a finite set containing an extremal projection cycle-vector
Above we study some structural properties of optimal projection cycles on certain finite sets. Below, we will use this properties to prove that there is a projection cycle on an explicitly given finite set that gives the supremum (2.4).
@@
Let $M_{i},1\leq i\leq 8$ be a maximum point of the function $g_{i}$ on the domain $\Omega=[0,1]^{3}$. And $F_{i},1\leq i\leq 6$ be a maximum point of the function $h_{i}$ on the face it’s defined.
@@
$U=\{T_{i},1\leq i\leq 8\}\cup\{M_{i},1\leq i\leq 8\}\cup\{F_{i},1\leq i\leq 6\}$
and let $U_{x},U_{y},U_{z}$ be the set of projections of elements of $U$ to the coordinate lines $Ox,Oy,Oz$ respectively. We will see that, the choice of $M_{i}$ and $F_{i}$ among the maximum points of $g_{i}$ and $h_{i}$ will not change the correctness of the following theorem.
@@
[…]
@@
$\max_{(p,\lambda)\in X_{S}}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}=\max_{(p,\lambda)\in X_{U_{x}\times U_{y}\times U_{z}}}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}$ (2.12)
@@
Let, $X^{\prime\prime}_{S}$ be the set of projection cycle-vectors $(p,\lambda)\in X_{S}$ that satisfies the propositions $(b),(c),(d)$ of the Theorem 2.2.
It is clear that, $X^{\prime}_{S}\subset X^{\prime\prime}_{S}$ and hence from the Lemma 2.5 we have,
$\max_{(p,\lambda)\in X_{S}}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}=\max_{(p,\lambda)\in X^{\prime\prime}_{S}}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}.$
@@
$\max_{(p,\lambda)\in X^{\prime\prime}_{S}}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}=\max_{(p,\lambda)\in X_{U_{x}\times U_{y}\times U_{z}}}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}$
Assume the contrary. Since, $U\subset S$ and $S=S_{x}\times S_{y}\times S_{z}$ we have $U_{x}\subset S_{x}$, $U_{y}\subset S_{y}$ and $U_{z}\subset S_{z}$ which imply $U_{x}\times U_{y}\times U_{z}\subset S$ and therefore we can write the inequality
$\max_{(p,\lambda)\in X^{\prime\prime}_{S}}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}\geq\max_{(p,\lambda)\in X_{U_{x}\times U_{y}\times U_{z}}}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}.$
Hence, there exist a projection cycle $(p^{0},\lambda^{0})\in X^{\prime\prime}_{S}$ such that
$\frac{\sum\lambda^{0}_{i}f(x^{0}_{i})}{\sum|\lambda^{0}_{i}|}>\max_{(p,\lambda)\in X_{U_{x}\times U_{y}\times U_{z}}}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}.$ (2.13)
Clearly, $(p^{0},\lambda^{0})\not\in X_{U_{x}\times U_{y}\times U_{z}}$.
If all points in $p^{0}$ with positive weight belong to $U$, then the set of the projections of them into the edges of the unit cube (which are all the possible locations of points of $p^{0}$ with negative weights) would belong to the set $U_{x}\times U_{y}\times U_{z}$ and hence we would have $p^{0}\subset U_{x}\times U_{y}\times U_{z}$.
Thus, there exist a point $x\in p^{0}$ with positive weight($\lambda^{0}(x)>0$) such that, $x\not\in U$. Moreover, let $(p^{0},\lambda^{0})$ be chosen to minimise the number of positive weighted points not belonging to the set $U$ among all projection cycle-vectors $(p,\lambda)\in X^{\prime\prime}_{S}$ satisfying (2.13).
By and Lemma 2.5 and Theorem 2.2 part $(d)$, we have that $x$ cannot be on the edges of the unit cube. Thus, we have the following cases:
@@
Consider the planes $x=\xi$ and $y=\gamma$. In each of these planes, at least one point of $(p^{0},\lambda^{0})$ with negative weight should exist. By Theorem 2.2, part $(c)$, these negative weighted points could only be on the edges, not adjacent to any of the vertices $(0,0,0)$ and $(1,1,1)$. The intersections of the planes $x=\xi$ and $y=\gamma$ and the edges are the points $(\xi,1,0),(\xi,0,1)$ and $(1,\gamma,0),(0,\gamma,1)$ respectively.
Since both of $(\xi,1,0)$ and $(1,\gamma,0)$ are well-ordered with $x$ and lying on an interior plane, by the Theorem 2.2, part $(b)$, none of $(\xi,1,0)$ and $(1,\gamma,0)$ can have a negative weight. Thus, both $(\xi,0,1)$ and $(0,\gamma,1)$ have negative weights. Moreover, since no other negative weighted points can exist on the planes $x=\xi$ and $y=\gamma$, we have:
@@
and thus, $\min(|\lambda^{0}(\xi,0,1)|,|\lambda^{0}(0,\gamma,1)|,|\lambda^{0}(x)|)=\lambda^{0}(x)$.
Let’s move the weights on the points $x$,$(\xi,0,1)$ and $(0,\gamma,1)$ to the points $F_{1}=(\xi_{1},\gamma_{1},0)$ and its two corresponding projections, namely $(\xi_{1},0,1)$ and $(0,\gamma_{1},1)$. Or in other words, we define
@@
$\lambda_{\varepsilon}=(-\varepsilon,\varepsilon,\varepsilon,\varepsilon,-\varepsilon,-\varepsilon)$
where, $\varepsilon=\lambda^{0}(x)$. Let $(p^{1},\lambda^{1})=(p^{0},\lambda^{0})+(p^{\prime},\lambda_{\varepsilon})$. Since, at each of the first three points of $p^{\prime}$, the weights defined in $(p^{\prime},\lambda_{\varepsilon})$ and $(p^{0},\lambda^{0})$ have opposite signs and
@@
by the Lemma 2.2 we have
$\sum|\lambda^{1}_{i}|\leq\sum|\lambda^{0}_{i}|$
@@
$=f(\xi_{1},\gamma_{1},0)-f(\xi_{1},1,0)-f(1,\gamma_{1},0)-f(\xi,\gamma,0)+f(\xi,1,0)+f(1,\gamma,0)=h(\xi_{1},\gamma_{1},0)-h(\xi,\gamma,0)\geq 0$
where the inequality follows from the choice of $F_{1}=(\xi_{1},\gamma_{1},0)$ as maximum of the function $h_{1}$. Thus, we get
$\frac{\sum\lambda^{1}_{i}f(x_{i})}{\sum|\lambda^{1}_{i}|}\geq\frac{\sum\lambda^{0}_{i}f(x_{i})}{\sum|\lambda^{0}_{i}|}>\max_{(p,\lambda)\in X_{U_{x}\times U_{y}\times U_{z}}}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}$
Clearly, $(p^{1},\lambda^{1})\in X^{\prime\prime}_{S}$ as no negative weighted point became positive weighted and vice versa. Moreover, $\lambda^{1}(x)=\lambda^{0}(x)-\lambda_{\varepsilon}(x)=0$ and hence, $(p^{1},\lambda^{1})$ has less number of positive weighted points not belonging $U$ than $(p^{0},\lambda^{0})$ has. Thus, we get a contradiction to the choice of $(p_{0},\lambda_{0})$.
@@
Let $x=(\xi,\gamma,\zeta)$. Consider the planes $x=\xi$, $y=\gamma$, $z=\zeta$. Their intersections with the edges of the unit cube that are not adjacent to the vertices not adjacent to any of the vertices $(0,0,0)$ and $(1,1,1)$ will be the points $(\xi,1,0),(\xi,0,1)$ and $(1,\gamma,0),(0,\gamma,1)$ and $(1,0,\zeta),(0,1,\zeta)$ respectively. To make the sum over the corresponding plane zero, at least one (maybe both) of two points in each plane should belong to $(p^{0},\lambda^{0})$ with a negative weight.
@@
For each $G_{i}$ we proceed by moving the weights at point $x$ and elements of $G_{i}$ to $M_{i}$ and its corresponding projections on the edges. For example, if all elements of $G_{1}$ belong to $p^{0}$, then we construct $p^{\prime}=(x,(\xi,0,1),(1,\gamma,0),(0,1,\zeta),(\xi_{1},\gamma_{1},\zeta_{1}),(\xi_{1},0,1),(1,\gamma_{!},0),(0,1,\zeta_{1}))$ and vector $\lambda_{\varepsilon}=(-\varepsilon,\varepsilon,\varepsilon,\varepsilon,\varepsilon,-\varepsilon,-\varepsilon,-\varepsilon)$ where $M_{1}=(\xi_{1},\gamma_{1},\zeta_{1})$ is the maximum point of $g_{1}$ defined above and $\varepsilon>0$.
As in Case 1, let $(p^{1},\lambda^{1})=(p^{0},\lambda^{0})+(p^{\prime},\lambda_{\varepsilon}).$ Since, at each of the first four points of $(p^{\prime},\lambda_{\varepsilon})$, the weights defined in $(p^{\prime},\lambda_{\varepsilon})$ and $(p^{0},\lambda^{0})$ have opposite signs, by the Lemma 2.2 for $\varepsilon=\min(\lambda)$ we have
@@
$\frac{\sum\lambda^{1}_{i}f(x_{i})}{\sum|\lambda^{1}_{i}|}\geq\frac{\sum\lambda^{0}_{i}f(x_{i})}{\sum|\lambda^{0}_{i}|}$
As in the first case the last inequality contradicts our assumption of $(p^{0},\lambda^{0})$. Similarly for all other sets we will get a contradiction.
Since all cases lead to the contradiction, the theorem is proved.
The proof implies that, the choice of $M_{i}$ and $F_{i}$ among the maximum points of $g_{i}$ and $h_{i}$ does not affect to the correctness of the theorem.
Now we are ready to prove the Theorem $2.1$.
Proof (Theorem 2.1).
Applying the Theorem 2.3 in the relation (2.5) we get
$E(f,\Omega)=\sup_{A,B,C}\left({\max_{(p,\lambda)\in X_{A\times B\times C}}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}}\right)=\sup_{A,B,C}\left({\max_{(p,\lambda)\in X_{(U_{x}\cup A)\times(U_{y}\cup B)\times(U_{z}\cup C)}}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}}\right)=$
$\sup_{A,B,C}\left(\max_{(p,\lambda)\in X_{U_{x}\times U_{y}\times U_{z}}}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}\right)=\max_{(p,\lambda)\in X_{U_{x}\times U_{y}\times U_{z}}}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}.$
@@
The proofs of the Theorem 2.3 and the Theorem 2.1 shows that, there exists a projection cycle-vector $(p,\lambda)\in X_{U_{x}\times U_{y}\times U_{z}}$ such that $(p,\lambda)$ satisfies all the propositions of the Theorem 2.2 and
@@
Moreover, it can be reconstructed by only knowing the weights in the vertices and all positive weighted points.

### v2

@@ 2.1 Optimal Projection Cycle-Vectors on a Finite Set
For any set $S=S_{x}\times S_{y}\times S_{z}$, where $S_{x},S_{y},S_{z}\subset[0,1]$ are some finite sets with $\{0,1\}\subset S_{x}\cap S_{y}\cap S_{z}$ and for any function $f\in C(\Omega)$ satisfying the inequalities (2.1), the intersection $S]^{(a)-(e)}\cap[S]_{f}$ is non-empty.
@@
$f_{n}(x,y,z)=f(x,y,z)+\frac{1}{n}(xy+xz+yz),\ \forall n\in\mathbb{N}.$
@@
$\max_{(p,\lambda)\in[S]}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}=\max_{(p,\lambda)\in[S]_{min}}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}=\max_{(p,\lambda)\in[S]_{min}}\frac{\sum\lambda_{i}(\lim_{n\rightarrow\infty}f_{n}(x_{i}))}{\sum|\lambda_{i}|}.$
Here, we can change the order of the limit and the maximum, as $S]_{min}$ is a finite set by Lemma 2.3 (considering vectors up to multiplication by a constant), and hence,
$\max_{(p,\lambda)\in[S]_{min}}\frac{\sum\lambda_{i}(\lim_{n\rightarrow\infty}f_{n}(x_{i}))}{\sum|\lambda_{i}|}=\lim_{n\rightarrow\infty}\max_{(p,\lambda)\in[S]_{min}}\frac{\sum\lambda_{i}f_{n}(x_{i})}{\sum|\lambda_{i}|}=\lim_{n\rightarrow\infty}\max_{(p,\lambda)\in[S]}\frac{\sum\lambda_{i}f_{n}(x_{i})}{\sum|\lambda_{i}|}.$
@@
$\lim_{n\rightarrow\infty}\max_{(p,\lambda)\in[S]}\frac{\sum\lambda_{i}f_{n}(x_{i})}{\sum|\lambda_{i}|}=\lim_{n\to\infty}\max_{(p,\lambda)\in[S]^{(a)-(e)}}\frac{\sum\lambda_{i}f_{n}(x_{i})}{\sum|\lambda_{i}|}=\lim_{n\to\infty}\max_{(p,\lambda)\in[S]^{(a)-(e)}_{min}}\frac{\sum\lambda_{i}f_{n}(x_{i})}{\sum|\lambda_{i}|}.$
@@
$\lim_{n\rightarrow\infty}\max_{(p,\lambda)\in[S]^{(a)-(e)}_{min}}\frac{\sum\lambda_{i}f_{n}(x_{i})}{\sum|\lambda_{i}|}=\max_{(p,\lambda)\in[S]^{(a)-(e)}_{min}}\frac{\sum\lambda_{i}(\lim_{n\rightarrow\infty}f_{n}(x_{i}))}{\sum|\lambda_{i}|}=\max_{(p,\lambda)\in[S]^{(a)-(e)}}\frac{\sum\lambda_{i}(f(x_{i}))}{\sum|\lambda_{i}|}.$
Now let’s choose $(p^{0},\lambda^{0})\in[S]^{(a)-(e)}$ such that,
$\frac{\sum\lambda^{0}_{i}(f(x_{i}))}{\sum|\lambda^{0}_{i}|}=\max_{(p,\lambda)\in[S]^{(a)-(e)}}\frac{\sum\lambda_{i}(f(x_{i}))}{\sum|\lambda_{i}|}.$
@@
$\frac{\sum\lambda^{0}_{i}(f(x_{i}))}{\sum|\lambda^{0}_{i}|}=\max_{(p,\lambda)\in[S]}\frac{\sum\lambda_{i}(f(x_{i}))}{\sum|\lambda_{i}|},$
and thus, $(p^{0},\lambda^{0})\in[S]_{f}$ by the definition of $S]_{f}$. Hence, $(p^{0},\lambda^{0})\in[S]^{(a)-(e)}\cap[S]_{f}$. ∎
@@ 2.2 Construction of a finite set containing an extremal projection cycle-vector
Above, we study some structural properties of optimal projection cycles on certain finite sets. Below, we will use these properties to prove that there is a projection cycle on an explicitly given finite set that gives the supremum (2.3).
Let $f:\Omega\rightarrow\mathbb{R}$ be a given continuous function of three variables satisfying conditions (2.1).
@@
Let $M_{i},1\leq i\leq 8$ be a maximum point of the function $g_{i}$ on the domain $\Omega=[0,1]^{3}$. And $F_{i},1\leq i\leq 6$ be a maximum point of the function $h_{i}$ on the face it is defined.
@@
$U=U(f)=\{T_{i},1\leq i\leq 8\}\cup\{M_{i},1\leq i\leq 8\}\cup\{F_{i},1\leq i\leq 6\},$
and let $U_{x},U_{y},U_{z}$ be the set of projections of elements of $U$ to the coordinate lines $Ox,Oy,Oz$ respectively. We will see that the choice of $M_{i}$ and $F_{i}$ among the maximum points of $g_{i}$ and $h_{i}$ will not change the correctness of the following theorem.
@@
Some of the points $T_{i},\ 1\leq i\leq 8$, $M_{j},\ 1\leq j\leq 8$, $F_{k},\ 1\leq k\leq 6$ may coincide. In this case, the set $U$ will narrow.
@@
$\max_{(p,\lambda)\in[S]}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}=\max_{(p,\lambda)\in[U_{x}\times U_{y}\times U_{z}]}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}.$ (2.11)
@@
Let, $S]^{(a)-(d)}$ be the set of projection cycle-vectors $(p,\lambda)\in[S]$ that satisfies propositions $(a),(b),(c),(d)$ of the Theorem 2.2.
It is clear that $S]^{(a)-(e)}\subset[S]^{(a)-(d)}$, and hence from Lemma 2.5 we have
$\max_{(p,\lambda)\in[S]}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}=\max_{(p,\lambda)\in[S]^{(a)-(d)}}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}.$
@@
$\max_{(p,\lambda)\in[S]^{(a)-(d)}}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}=\max_{(p,\lambda)\in[U_{x}\times U_{y}\times U_{z}]}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}.$
Since $U\subset S$ and $S=S_{x}\times S_{y}\times S_{z}$ we have $U_{x}\subset S_{x}$, $U_{y}\subset S_{y}$ and $U_{z}\subset S_{z}$ which imply $U_{x}\times U_{y}\times U_{z}\subset S$ and therefore we can write the inequality
$\max_{(p,\lambda)\in[S]^{(a)-(d)}}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}\geq\max_{(p,\lambda)\in[U_{x}\times U_{y}\times U_{z}]}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}.$
Assume the contrary to the theorem, let there exists a projection cycle $(p^{0},\lambda^{0})\in[S]^{(a)-(d)}$ such that
$\frac{\sum\lambda^{0}_{i}f(x^{0}_{i})}{\sum|\lambda^{0}_{i}|}>\max_{(p,\lambda)\in[U_{x}\times U_{y}\times U_{z}]}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}.$ (2.12)
Clearly, $(p^{0},\lambda^{0})\not\in[U_{x}\times U_{y}\times U_{z}]$. Moreover we can choose it to be optimal in the set $S$ for the function $f$, that is, $(p^{0},\lambda^{0})\in[S]_{f}$.
If all points in $p^{0}$ with positive weight belong to $U$, then the set of their projections on the edges of the unit cube (which are all the possible locations of points of $p^{0}$ with negative weights) would belong to the set $U_{x}\times U_{y}\times U_{z}$, and hence we would have $p^{0}\subset U_{x}\times U_{y}\times U_{z}$.
[…]
We define our moves based on the location of $x$ in the unit cube. Thus, we have the following cases:
@@
Consider the planes $x=\xi$ and $y=\gamma$. In each of these planes, there should be at least one point $(p^{0},\lambda^{0})$ with negative weight. By Theorem 2.2, part $(c)$, these negative weighted points could only be on the edges, not adjacent to any of the vertices $(0,0,0)$ and $(1,1,1)$. The intersections of the planes $x=\xi$ and $y=\gamma$ and the edges are the points $(\xi,1,0),(\xi,0,1)$ and $(1,\gamma,0),(0,\gamma,1)$, respectively.
Since both $(\xi,1,0)$ and $(1,\gamma,0)$ are well-ordered with $x$ and lie on an interior plane, by Theorem 2.2, part $(b)$, none of $(\xi,1,0)$ and $(1,\gamma,0)$ can have a negative weight. Thus, both $(\xi,0,1)$ and $(0,\gamma,1)$ have negative weights. Moreover, since no other negative weighted points can exist on the planes $x=\xi$ and $y=\gamma$, we have:
@@
and thus $\min(|\lambda^{0}(\xi,0,1)|,|\lambda^{0}(0,\gamma,1)|,|\lambda^{0}(x)|)=\lambda^{0}(x)$.
[…]
@@
$\lambda_{\varepsilon}=(-\varepsilon,\varepsilon,\varepsilon,\varepsilon,-\varepsilon,-\varepsilon),$
where $\varepsilon=\lambda^{0}(x)$. Let $(p^{1},\lambda^{1})=(p^{0},\lambda^{0})+(p^{\prime},\lambda_{\varepsilon})$. Since, at each of the first three points of $p^{\prime}$, the weights defined in $(p^{\prime},\lambda_{\varepsilon})$ and $(p^{0},\lambda^{0})$ have opposite signs and
@@
by Lemma 2.2 we have
$\sum|\lambda^{1}_{i}|\leq\sum|\lambda^{0}_{i}|.$
@@
$f(\xi_{1},\gamma_{1},0)-f(\xi_{1},1,0)-f(1,\gamma_{1},0)-f(\xi,\gamma,0)+f(\xi,1,0)+f(1,\gamma,0)=h(\xi_{1},\gamma_{1},0)-h(\xi,\gamma,0)\geq 0,$
where the inequality follows from the choice of $F_{1}=(\xi_{1},\gamma_{1},0)$ as the maximum of the function $h_{1}$. Thus, we get
$\frac{\sum\lambda^{1}_{i}f(x_{i})}{\sum|\lambda^{1}_{i}|}\geq\frac{\sum\lambda^{0}_{i}f(x_{i})}{\sum|\lambda^{0}_{i}|}>\max_{(p,\lambda)\in[U_{x}\times U_{y}\times U_{z}]}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}.$
This proves that, $(p^{1},\lambda^{1})\in[S]_{f}$. This makes the weight at point $x$ equal to zero in the new projection cycle. A similar movement of the weights can be constructed if $x$ lies on one of the other faces too.
@@
Let $x=(\xi,\gamma,\zeta)$. Consider the planes $x=\xi$, $y=\gamma$, $z=\zeta$. Their intersections with the edges of the unit cube that are not adjacent to the vertices not adjacent to any of the vertices $(0,0,0)$ and $(1,1,1)$ will be the points $(\xi,1,0),(\xi,0,1)$ and $(1,\gamma,0),(0,\gamma,1)$ and $(1,0,\zeta),(0,1,\zeta)$, respectively. To make the sum over the corresponding plane zero, at least one (maybe both) of two points in each plane should belong to $(p^{0},\lambda^{0})$ with a negative weight.
@@
For each $G_{i}$ we proceed by moving the weights at point $x$ and elements of $G_{i}$ to $M_{i}$ and its corresponding projections on the edges. For example, if all elements of $G_{1}$ belong to $p^{0}$, then we construct $p^{\prime}=(x,(\xi,0,1),(1,\gamma,0),(0,1,\zeta),(\xi_{1},\gamma_{1},\zeta_{1}),(\xi_{1},0,1),(1,\gamma_{1},0),(0,1,\zeta_{1}))$ and vector $\lambda_{\varepsilon}=(-\varepsilon,\varepsilon,\varepsilon,\varepsilon,\varepsilon,-\varepsilon,-\varepsilon,-\varepsilon)$ where $M_{1}=(\xi_{1},\gamma_{1},\zeta_{1})$ is the maximum point of $g_{1}$ defined above and $\varepsilon>0$.
As in Case 1, let $(p^{1},\lambda^{1})=(p^{0},\lambda^{0})+(p^{\prime},\lambda_{\varepsilon}).$ Since, at each of the first four points of $(p^{\prime},\lambda_{\varepsilon})$, the weights defined in $(p^{\prime},\lambda_{\varepsilon})$ and $(p^{0},\lambda^{0})$ have opposite signs, by Lemma 2.2 for
$\varepsilon=\min(|\lambda(x)|,|\lambda((\xi,0,1))|,|\lambda((1,\gamma,0))|,|\lambda((0,1,\zeta))|$
we have
@@
$\frac{\sum\lambda^{1}_{i}f(x_{i})}{\sum|\lambda^{1}_{i}|}\geq\frac{\sum\lambda^{0}_{i}f(x_{i})}{\sum|\lambda^{0}_{i}|}.$
Again, this proves that, $(p^{1},\lambda^{1})\in[S]_{f}$. On the other hand, choice of $\varepsilon$ makes one of the weights of the points $x,(\xi,0,1),(1,\gamma,0),(0,1,\zeta)$ to have zero weight in $(p^{1},\lambda^{1})$. Hence, after a finite number of copies of the above change for some sets $G_{i}$, the weight in $x$ will be equal to zero in the resulting projection cycle-vector.
Thus, we can proceed a finite number of described moves and obtain a projection cycle-vector from $U_{x}\times U_{y}\times U_{z}]$, which is optimal in $S$. ∎
The proof implies that the choice of $M_{i}$ and $F_{i}$ among the maximum points of $g_{i}$ and $h_{i}$ does not affect the correctness of the theorem. Now, we are ready to state our main result.
Theorem 2.3.
Let $f\in C(\Omega)$ be a function satisfying the relations (2.1). Then
$E(f,\Omega)=\max_{(p,\lambda)\in[U_{x}\times U_{y}\times U_{z}]^{(a)-(e)}_{min}}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|},$ (2.13)
where $U_{x},U_{y},U_{z}$ are the projections of $U$ to the coordinate lines $Ox,Oy,Oz$ respectively.
Proof.
Applying Theorem 2.2 to the relation (2.4) we get
$\begin{split}E(f,\Omega)=\sup_{A,B,C}\left({\max_{(p,\lambda)\in[A\times B\times C]}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}}\right)=\\ =\sup_{A,B,C}\left({\max_{(p,\lambda)\in[(U_{x}\cup A)\times(U_{y}\cup B)\times(U_{z}\cup C)]}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}}\right)=\\ =\sup_{A,B,C}\left(\max_{(p,\lambda)\in[U_{x}\times U_{y}\times U_{z}]}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}\right)=\max_{(p,\lambda)\in[U_{x}\times U_{y}\times U_{z}]}\frac{\sum\lambda_{i}f(x_{i})}{\sum|\lambda_{i}|}.\end{split}$ (2.14)
Hence, we are finished by the Lemma 2.4 and Lemma 2.5. ∎
@@
The above-mentioned papers [11],[12], [20] give similar theorems. Their set corresponding to $U$ in our case, was the corners of the unit square in papers [11] and [12], and points from a predefined grid in [20]. While all these are independent of $f$, $U$ depends on the function $f$ in the above theorem.
Remark 2.3.
The Theorem 2.3 shows that there exists a minimal projection cycle-vector $(p,\lambda)\in[U_{x}\times U_{y}\times U_{z}]$ such that $(p,\lambda)$ satisfies all the propositions of Theorem 2.2 and
@@
In the Appendix, we have shown that there are at most 123 such projection cycle-vectors in the set $U$. Moreover, any of them can be reconstructed only by knowing the weights in the vertices and all the positive weighted points(if two of those points coincide, then thinking them as separate and adding their weights after the weight calculations are done) by using the definitions of $M_{i}$ and $F_{i}$. For example, if $M_{i}$ has weight $\lambda$, we will add $\lambda$ to all of the points on $G_{i}$.